OpenStudy (anonymous):

9x^4=25x^2-16 solve using an appropriate substitution

jimthompson5910 (jim_thompson5910):

Let z = x^2 So z^2 = (x^2)^2 = x^4 This means 9x^4=25x^2-16 becomes 9z^2=25z-16

jimthompson5910 (jim_thompson5910):

Solve that last equation for z. Then use the solutions in terms of z to find the solutions in terms of x.

OpenStudy (anonymous):

okay but heres where im stuck: 9z^2-25y+16=0.. how do i factor that?

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to 9*16 = 144 AND add to -25

OpenStudy (anonymous):

multiply 9 and 16?

jimthompson5910 (jim_thompson5910):

yes that gives you 144

jimthompson5910 (jim_thompson5910):

you need to find two numbers that multiply to 144 and add to -25

OpenStudy (anonymous):

i thought we were finding two numbers that multiplied to equal 16 and which add to -25?

jimthompson5910 (jim_thompson5910):

that only works if the leading coefficient is 1

OpenStudy (anonymous):

oh really now

OpenStudy (anonymous):

so in any other case you are multiplying the first and last part?

jimthompson5910 (jim_thompson5910):

yes given ax^2+bx+c, it factors when you can find two numbers that multiply to ac and add to b

jimthompson5910 (jim_thompson5910):

example: Factor x^2+5x+6 that's really 1x^2+5x+6 So we need to find two numbers that multiply to 1*6 = 6 and add to 5...those numbers are 2 and 3 So it factors to (x+2)(x+3) Note: this last jump only works if the leading coefficient is 1

OpenStudy (anonymous):

ok so if it was 2x^2+5x+6 youd be looking for what times what equals 12 and adds to 5?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

okay so its gonna be -9 and -16

jimthompson5910 (jim_thompson5910):

good, now you can replace -25z with -9z-16z to go from 9z^2-25z+16=0 to 9z^2-9z-16z+16=0

jimthompson5910 (jim_thompson5910):

Then you factor by grouping

OpenStudy (anonymous):

(9y-16)(y-1)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

now i need to convert back to x's?

OpenStudy (anonymous):

nailed it. nice

jimthompson5910 (jim_thompson5910):

you can if you want, doing so will turn (9y-16)(y-1) into (9x^2-16)(x^2-1) assuming you let y = x^2

jimthompson5910 (jim_thompson5910):

Then you can factor each piece separately

OpenStudy (anonymous):

ya they expect me to. my answer matches the books now

jimthompson5910 (jim_thompson5910):

alright great

OpenStudy (anonymous):

perfect. thanks so much for the help

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

that little detail about a*c was a big break through for me

jimthompson5910 (jim_thompson5910):

I'm glad I could shed light on it